If a young man rides his motorcycle at 25 km/hr, he had to spend Rs.2 per km on petrol. If he
rides at a faster s peed of 40 km/hr, the petrol cost increases at Rs.5 per km. He has Rs.100 to spend on petrol and wishes to find what is the maximum distance he can travel within one hour.Express this as LPP and solve it graphically

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
From given datayoung man rides his motorcycle at 25 km/hr, he had to spend Rs.2 per km on petrol. If he rides at a faster s peed of 40 km/hr, the petrol cost increases at Rs.5 per km. He has Rs.100 to spend on petrol
By writing the equations we get
2x+5y\leq100
(x/25)+(y/40)\leq1
Z=x+y is to be maximized
We get maximum at x=50/3 and y=40/3 so maximum distance is 30kms